# ThyroidEffi 1.0: A Cost-Effective System for High-Performance Multi-Class Thyroid Carcinoma Classification

### v1

@@ 4.4 Real-World Application Challenges
Real-world deployment observations indicate a tendency of the model to misclassify Bethesda V (Suspicious for Malignancy) cases as Bethesda VI (Malignant). This issue likely arises from the inherent diagnostic ambiguity between these two categories, even for experienced cytopathologists. Furthermore, potential variations in data distribution exist between the 108 Military Central Hospital dataset (used for training and initial testing) and the Hung Viet Hospital dataset (used for external validation). These distribution shifts may stem not only from differences in patient populations and imaging protocols between the facilities but also from temporal drifts in data characteristics. These findings underscore the critical need for continuous monitoring and periodic retraining of the model using real-world data to enhance its performance and robustness in diverse clinical environments over time.
@@ 6 Summary table
  • Binary Classification Limitation: Many studies focus on binary classification (benign/malignant), neglecting the full multi-class categorization of the Bethesda System for Reporting Thyroid Cytopathology (TBSRTC).
@@ A.1 Study Population and Inclusion/Exclusion Criteria
  • Availability of sufficient cytological material in FNAB samples to allow classification according to the Bethesda System (categorized as B2, B5, or B6).
@@ A.6 Data Collection Process and Quality Considerations
  • Each FNAB image was associated with a unique patient identifier and labeled according to the Bethesda System for Reporting Thyroid Cytopathology (B2, B5, or B6) by expert cytopathologists.
@@ B.2 Gold Standard: Ground Truthing Process
FNAB samples were classified according to the Bethesda System for Reporting Thyroid Cytopathology (TBSRTC) guidelines by 3 board-certified cytopathologists with an average of 8 years of experience. Each sample was independently reviewed by at least two cytopathologists. Discrepancies in classification were resolved through discussion and consensus. In cases where consensus could not be reached, a third expert was consulted. Image quality was controlled by ensuring all images met pre-defined criteria for resolution, brightness, and contrast. Any images failing to meet these criteria were excluded from the study.
@@ C.4 Data Imbalance Analysis and Adjustment
  • Benign (B2): 482 images (26.7%)
  • Suspicious for Malignancy (B5): 541 images (30.0%)
  • Malignant (B6): 871 images (48.3%)
As can be seen, the Malignant (B6) category had considerably more images than the other two categories. This imbalance could potentially bias the model towards the majority class.
@@ D.2 Model Output
The model outputs a three-dimensional vector of probabilities, representing the predicted likelihood for each of the three Bethesda categories: Benign (B2), Suspicious for Malignancy (B5), and Malignant (B6). Specifically, for a given input image, the model produces the following output:
$\mathbf{p}=[p_{B2},p_{B5},p_{B6}]$
@@
  • $p_{B2}$ represents the probability that the image is Benign (B2).
  • $p_{B5}$ represents the probability that the image is Suspicious for Malignancy (B5).
  • $p_{B6}$ represents the probability that the image is Malignant (B6).
@@
where $i\in\{B2,B5,B6\}$.
@@ D.4 Data Augmentation Technique: YOLOv10 Rationale and Overfitting Considerations
Our training protocol is built on curriculum learning principles by processing image subsets in a defined sequence: E, D, C, B, and A. In this sequence, sets E, D, and C consist of small cropped regions from the main image that capture localized, high-noise features, while sets B and A encompass larger portions or the entire image, providing global, patient-level context. This ordering leverages the observation that more recent training examples tend to have a stronger influence on model updates. By first challenging the model with localized, noisy regions and then progressively introducing broader, cleaner images, we aim to foster the learning of both fine-grained details and comprehensive contextual features.
Although we have not yet benchmarked this sequential strategy against a random ordering baseline, our approach is firmly grounded in established curriculum learning concepts. The gradual transition from difficult local sub-regions to more informative global views is intended to enhance the retention of newly acquired features, thereby improving multi-scale feature extraction and overall diagnostic performance.
@@ D.8 Multi-scale and Multi-region Analysis: Transformer-based Approach
    Final Label Prediction: The token corresponding to the primary region (typically $T_{0}$) is passed through a feed-forward neural network (FFNN) to produce the final three-dimensional output vector, representing the predicted class labels for the entire image.
@@ E.5 External Validation Set Characteristics and Comparison to Training Data
The distribution of Bethesda categories in the external set was as follows: B2: [300], B5: [315], B6: [400].
@@ F.3 System Adoption and Deployment
The ThyroidEffi Basic model has been deployed at Hung Viet Hospital since January 2025. It is integrated into the clinical workflow as follows: After a cytopathologist reviews a slide, they upload the digital image to the ThyroidEffi system via a secure web interface. The system processes the image and returns a classification (B2, B5, B6) along with Grad-CAM visualizations. This output is then displayed alongside the original image in the cytopathologist’s workstation.
@@
Initial feedback from the cytopathologists has been positive. They report that the system’s visualizations help them to identify key diagnostic features more quickly and confidently, particularly in challenging cases. One cytopathologist noted, "The system has helped me catch a couple of subtle malignant cases that I might have otherwise missed." They have also noted that the system is particularly helpful in distinguishing between B5 (Suspicious) and B6 (Malignant) cases, which are often difficult to differentiate.

### v2

@@ 4.4 Real-World Application Challenges
Real-world deployment and external validation observations revealed a performance disparity compared to initial testing. Specifically, the model demonstrated a tendency to misclassify Indeterminate/Suspicious cases, sometimes assigning them to the Malignant category, which contributed to a lower overall performance on the external validation set. While both the training/initial testing dataset from 108 Military Central Hospital and the external validation dataset from Hung Viet Hospital represent real-world data, variations in performance between institutions are a known challenge in medical imaging due to differences in data distribution.
Potential causes for these observed performance differences and the model’s generalizability challenges likely include:
  • Patient Demographics: Differences in patient populations, prevalence of specific thyroid pathologies, or demographic factors between the two hospitals.
  • Imaging Protocols and Equipment: Variations in ultrasound equipment models, imaging settings, and FNAB acquisition protocols between the facilities can lead to distinct visual characteristics in the slide images.
  • Slide Preparation and Staining Variability: Differences in laboratory procedures for sample processing, slide preparation, and staining techniques can introduce variations in cellular morphology, background artifacts, and staining intensity.
  • Pathologist Variability: Despite standardized guidelines like TBSRTC, subtle differences in subjective interpretation and diagnostic thresholds among pathologists at different institutions can influence ground truth labeling. This is particularly pronounced for indeterminate categories, which are inherently more subjective within TBSRTC and prone to higher inter-observer variability, thus contributing to less consistent ground truth across institutions.
  • Temporal Data Drift: Changes in equipment calibration, protocols, or even disease characteristics over time could lead to the external validation data having subtly different feature distributions than the training data.
The lower performance observed on the external validation set indicates that the current model’s generalizability across different clinical sites and potentially over time presents a challenge that warrants further attention.
Future research will focus on further elucidating the specific contributions of these potential causes to model performance variability. We also plan to continuously improve the model’s robustness and generalizability through strategies such as continuous learning, which can help the model adapt to and capture changes in data distribution over time and across different clinical environments.
@@ 6 Summary table
  • Binary Classification Limitation: Many studies focus on binary classification (benign/malignant), neglecting the full multi-class categorization of TBSRTC.
@@
Glossary
AUC
Area Under the Curve
Benign
Bethesda II: Observation / Non-Surgical Management
CNN
Convolutional Neural Network
CODT
Correlation Optical Diffraction Tomography
FFNN
Feed-Forward Neural Network
FNAB
Fine Needle Aspiration Biopsy
Grad-CAM
Gradient-weighted Class Activation Mapping
Indeterminate/Suspicious
Bethesda I, III, IV, V: Further Investigation / Consideration for Intervention
IOPD
Intraoperative Thyroid Nodule Diagnosis
Malignant
Bethesda VI: Surgical Intervention
PCA
Principal Component Analysis
PTC
Papillary Thyroid Carcinoma
ROI
Region of Interest
SVM
Support Vector Machine
t-SNE
t-distributed Stochastic Neighbor Embedding
TBSRTC
The Bethesda System for Reporting Thyroid Cytopathology
UMAP
Uniform Manifold Approximation and Projection
WSI
Whole Slide Image
@@ A.1 Study Population and Inclusion/Exclusion Criteria
  • Availability of sufficient cytological material in FNAB samples to allow classification according to the Bethesda System (categorized as Benign, Indeterminate/Suspicious, or Malignant).
@@ A.6 Data Collection Process and Quality Considerations
  • Each FNAB image was associated with a unique patient identifier and labeled according to TBSRTC (Benign, Indeterminate/Suspicious, or Malignant) by expert cytopathologists.
@@ B.2 Gold Standard: Ground Truthing Process
FNAB samples were classified according to TBSRTC guidelines by 3 board-certified cytopathologists with an average of 8 years of experience. Each sample was independently reviewed by at least two cytopathologists. Discrepancies in classification were resolved through discussion and consensus. In cases where consensus could not be reached, a third expert was consulted. Image quality was controlled by ensuring all images met pre-defined criteria for resolution, brightness, and contrast. Any images failing to meet these criteria were excluded from the study.
@@ C.2 Missing Data Management
Note that images corresponding to Bethesda I samples were included in the Indeterminate/Suspicious category if they met the predefined quality criteria for cellularity and staining, while those deemed non-diagnostic due to inadequate material or poor quality were excluded as part of standard data preparation."
@@ C.4 Data Imbalance Analysis and Adjustment
As can be seen, Malignant category had considerably more images than the other two categories. This imbalance could potentially bias the model towards the majority class.
@@ D.2 Model Output
The model outputs a three-dimensional vector of probabilities, representing the predicted likelihood for each of the three Bethesda categories: Benign, Indeterminate/Suspicious, and Malignant. Specifically, for a given input image, the model produces the following output:
$\mathbf{p}=[p_{\text{Benign}},p_{\text{Indeterminate/Suspicious}},p_{\text{Malignant}}]$
@@
  • $p_{Benign}$ represents the probability for the Benign category.
  • $p_{\text{Indeterminate/Suspicious}}$ represents the probability for the Indeterminate/Suspicious category.
  • $p_{Malignant}$ represents the probability for the Malignant category.
@@
where $i\in\{Benign,Indeterminate/Suspicious,Malignant\}$.
@@ D.4 Data Augmentation Technique: YOLOv10 Rationale and Overfitting Considerations
Our training protocol incorporates principles of curriculum learning by processing image subsets in a defined sequence: E, D, C, B, and A. In this sequence, sets E, D, and C consist of smaller cropped regions from the main image that capture localized, high-noise features, while sets B and A encompass larger portions or the entire image, providing global, patient-level context. This ordering leverages the observation that more recent training examples tend to have a stronger influence on model updates, allowing us to strategically prioritize the learning of increasingly complex and contextual features. By first challenging the model with localized, noisy regions and then progressively introducing broader, cleaner images, we aim to foster the effective learning and integration of both fine-grained details and comprehensive contextual features across multiple scales.
This sequential approach is firmly grounded in established curriculum learning concepts, where presenting examples in a structured manner can enhance learning efficiency and feature retention. The gradual transition from difficult local sub-regions to more informative global views is intended to enhance the retention of newly acquired features from simpler examples while building upon them with more complex ones, thereby improving multi-scale feature extraction and overall diagnostic performance by facilitating the model’s ability to consolidate understanding across different levels of detail and context. This structured learning paradigm forms a core component of our novel training strategy.
@@ D.8 Multi-scale and Multi-region Analysis: Transformer-based Approach
    Final Label Prediction: The token corresponding to the primary region (typically $T_{0}$) is passed through a FFNN to produce the final three-dimensional output vector, representing the predicted class labels for the entire image.
@@ E.5 External Validation Set Characteristics and Comparison to Training Data
The distribution of Bethesda categories in the external set was as follows: Benign: [300], Indeterminate/Suspicious: [315], Malignant: [400].
@@ F.3 System Adoption and Deployment
The ThyroidEffi Basic model was deployed at Hung Viet Hospital in January 2025, after undergoing external validation in 1015 cases from July 2024 to November 25, 2024. It is integrated into the clinical workflow as follows: After a cytopathologist reviews a slide, they upload the digital image to the ThyroidEffi system via a secure web interface. The system processes the image and returns a classification (Benign, Indeterminate/Suspicious, Malignant) along with Grad-CAM visualizations. This output is then displayed alongside the original image in the cytopathologist’s workstation.
@@
Initial feedback from the cytopathologists has been positive. They report that the system’s visualizations help them to identify key diagnostic features more quickly and confidently, particularly in challenging cases. One cytopathologist noted, "The system has helped me catch a couple of subtle malignant cases that I might have otherwise missed." They have also noted that the system is particularly helpful in distinguishing between Indeterminate/Suspicious and Malignant cases, which are often difficult to differentiate.